### v1

# What Should Embeddings Embed? Autoregressive Models Represent Latent Generating Distributions

Abstract
Autoregressive language models have demonstrated a remarkable ability to extract latent structure from text. The embeddings from large language models have been shown to capture aspects of the syntax and semantics of language. But what should embeddings represent? We connect the autoregressive prediction objective to the idea of constructing predictive sufficient statistics to summarize the information contained in a sequence of observations, and use this connection to identify three settings where the optimal content of embeddings can be identified: independent identically distributed data, where the embedding should capture the sufficient statistics of the data; latent state models, where the embedding should encode the posterior distribution over states given the data; and discrete hypothesis spaces, where the embedding should reflect the posterior distribution over hypotheses given the data. We then conduct empirical probing studies to show that transformers encode these three kinds of latent generating distributions, and that they perform well in out-of-distribution cases and without token memorization in these settings.
1 Introduction
Autoregressive language models (LMs) are trained to predict the next token in a sequence (e.g., Bengio et al., 2000). Many large language models (LLMs) use the autoregressive objective for pretraining (e.g., Radford et al., 2019), and their document-level embeddings have been shown to capture elements of latent structure that appear in text, such as agent properties (Andreas, 2022) and syntax (Hewitt and Manning, 2019). However, it remains unclear why embeddings represent these elements, and, more formally, what information an embedding learned by a capable autoregressive predictor should capture. An understanding of what LLMs represent and why is important to the design and evaluation of architectures and optimizers, and to supporting transparency and safety.
The representations formed by LLMs are challenging to analyze partly due to polysemanticity, where a neuron may activate for several distinct concepts Cunningham et al. (2023). Previous work has developed methods for probing LLM internal representations for specific concepts (Li et al., 2023; Meng et al., 2022; Zheng et al., 2023; Tenney et al., 2019). These efforts can be guided more effectively by developing a general theory of what aspects of the data embeddings should represent.
In this work, we investigate several cases where the representations of autoregressive LMs can be formally connected with those of a Bayes-optimal agent. By linking the autoregressive objective to finding predictive sufficient statistics, we show that optimal content of embeddings can be identified in 1) independent identically distributed data, where the embedding should capture the sufficient statistics of the data; 2) latent state models, where the embedding should encode the posterior distribution over states given the data; and 3) discrete hypothesis spaces, where the embedding should reflect the posterior distribution over hypotheses given the data. We use probing methods to confirm that the relevant information can be decoded from LM embeddings. We also show that content that appears to be of a similar difficulty but not expected to be captured via predictive sufficient statistics is more challenging to recover from LM embeddings by probing.
Our analysis suggests that LLMs should represent latent structure that captures the posterior distribution over the generative process underlying text. Specifically, latent structure such that the next word $x_{n+1}$ is independent from previous words $x_{1:n}$ when conditioned on that structure should be recoverable from LLM embeddings. Additionally, by linking LM representations to Bayesian inference, our approach suggests that it may be effective to evaluate and interpret the behavior of LMs through comparison to more interpretable Bayes-optimal agents.
2 Related work
The embeddings produced by language models have been investigated in detail (Gupta et al., 2015; Köhn, 2015; Ettinger et al., 2016; Adi et al., 2017; Hupkes et al., 2018); for reviews, see Rogers et al. (2020) and Belinkov (2022). They have been shown to capture different aspects of the latent structure of text, including part of speech (Shi et al., 2016; Belinkov et al., 2017), syntactic number (Giulianelli et al., 2018; Conneau et al., 2018), sentence structure (Tenney et al., 2019; Hewitt and Manning, 2019; Liu et al., 2019; Lin et al., 2019), entity attributes (Gupta et al., 2015; Grand et al., 2022), sentiment (Radford et al., 2017), semantic roles (Ettinger et al., 2016; Tenney et al., 2019), world states (Li et al., 2021), and agent properties (Andreas, 2022). However, our work is motivated from a fundamentally different perspective. Instead of focusing on what is captured in the embeddings of these models, in this paper we explore why these particular kinds of structure might be represented as a consequence of statistical properties of the training data.
Several previous papers have analyzed LLMs by making a connection to Bayesian inference. Of these, Xie et al. (2021), McCoy et al. (2023), and Wang et al. (2024) analyze the in-context learning behavior of LLMs. However, we study what models should encode based on the autoregressive objective that is typically used to train LLMs. Zhang et al. (2023) and Zheng et al. (2023) also connect LLM embeddings to Bayesian inference, but they focus on topic models embedded in LLMs, while we extend the connection to more general cases.
Metalearned RNNs have also been shown to encode information equivalent to a Bayesian posterior distribution (Mikulik et al., 2020). Furthermore, recent work has also demonstrated that transformers behave like the Bayes-optimal predictor in linear regression settings (Panwar et al., 2024; Garg et al., 2022; Akyürek et al., 2023) and can approximate the posterior predictive distributions of probabilistic models such as Gaussian processes and Bayesian neural networks (Müller et al., 2022). We extend this analysis to general autoregressive language models and consider more general generative processes and what posterior distributions they should capture in these cases.
Figure 1: Three data generation processes where prediction of the next token $x_{n+1}$ is independent from previous tokens $x_{1:n}$ given a predictive sufficient statistic. The left corresponds to exchangeable data, the middle to latent state models, and the right to discrete hypotheses. The relevant predictive sufficient statistics are the sufficient statistic for $\theta$ (or $p(\theta|x_{1:n})$, $p(z_{n+1}|x_{1:n})$, and $p(h|x_{1:n})$ respectively). We show the embeddings learned by autoregressive transformers represent this information.
3 Three cases where optimal embeddings can be identified
Assume we have a sequence $x_{1:n}$ and an autoregressive language model (LM) that predicts the next item in the sequence, $p(x_{n+1}|x_{1:n})$. We denote the LM embedding for sequence $x_{1:n}$ as $\phi_{n}=f(x_{1:n})$. The distribution $p(x_{n+1}|x_{1:n})$ is some function $g(\phi_{n})$ of this embedding, with that function implemented by the final layers of the neural network instantiating the LM. That is, the probability of the next element in the sequence $x_{n+1}$ only depends on $\phi_{n}$. This establishes our basic question: what should $\phi_{n}$ represent in order to accurately predict $x_{n+1}$?
In this section we describe three cases where this question can be answered analytically (Figure 1). First, when $x_{1:n}$ are independently sampled conditioned on an unknown parameter, $\phi_{n}$ needs only represent the sufficient statistic of this sequence. Second, when $x_{1:n}$ are generated by a state space model (in the discrete case, a hidden Markov model), $\phi_{n}$ need only represent the posterior distribution over states given $x_{1:n}$. Finally, when $x_{1:n}$ are sampled independently from one of a discrete set of latent distributions, $\phi_{n}$ need only represent the posterior distribution over hypotheses about that distribution given $x_{1:n}$. In each case we explain how $p(x_{n+1}|x_{1:n})$ factorizes to make it possible for $x_{1:n}$ to be summarized by some $\phi_{n}$ and identify the form of the corresponding $g(\phi_{n})$.
3.1 General approach
The key idea behind our approach is that we can identify situations where the embedding $\phi_{n}$ contains all of the information from $x_{1:n}$ required to predict $x_{n+1}$. This idea is directly related to the notion of a sufficient statistic (Gelman et al., 2004). Given a distribution $p(x)$ with parameters $\theta$, a statistic $s(x)$ is sufficient for $\theta$ if the conditional distribution of $x$ given $s$ does not depend on $\theta$. In other words, if we only know $s$, we can estimate $\theta$ just as well as if we know the actual value of $x$. For example, if $x$ is a set of independently and identically distributed draws from a Gaussian distribution and $\theta$ is the mean of that distribution, then the mean of the sample $x$ is a sufficient statistic for $\theta$. The generative process can be rewritten to generate the sample mean $s$ given $\theta$ and then generate the variation around that sample mean independently from $\theta$, and $s$ contains all the information relevant to estimating $\theta$.
In the autoregressive setting, we care about predictive sufficiency (Bernardo and Smith, 2000). A statistic $s(x_{1:n})$ is predictive sufficient for the sequence $x_{1:n}$ if
$p(x_{n+1}|x_{1:n})=p(x_{n+1}|s(x_{1:n})).$ (1)
If a model performs autoregressive modeling perfectly, its embedding should represent a predictive sufficient statistic. Our three cases thus correspond to settings where predictive sufficient statistics can be easily identified and could plausibly be represented by a neural network.
3.2 Case 1: Exchangeable models
Predictive sufficiency is particularly straightforward to establish in exchangeable models, where the probability of a sequence remains the same under permutation of the order of its elements. That is, a sequence is exchangeable if $p(x_{1:N})=p(x_{\pi(1:N)})$ for some permutation $\pi$. Any exchangeable model can be re-expressed in terms of the $x_{i}$ being sampled independently and identically distributed according to a latent distribution $p(x|\theta)$ parameterized by $\theta$, with $p(x_{1:N})=\int_{\theta}\prod_{i}p(x_{i}|\theta)p(\theta)\,d\theta$ (Gelman et al., 2004). This idea leads to the following proposition:
Proposition. Given an exchangeable sequence $x_{1:N}$ where each $x_{i}$ is of dimension $d_{x}$, and given functions $f:\mathbb{R}^{nd_{x}}\mapsto\mathbb{R}^{d_{m}}$, $g:\mathbb{R}^{d_{m}}\mapsto\mathbb{R}^{d_{x}}$ such that, for each $1\leq n\leq N$, $g\circ f(x_{1:n})=p(x_{n+1}|x_{1:n})\;\forall x_{n+1}$, $f(x_{1:n})$ is a sufficient statistic for $x_{1:n}$.
In other words, if we have a perfect autoregressive predictor that is composable into $g\circ f$, the output of $f$ is a sufficient statistic for its sequence input.
Proof.
The result follows from the fact that for exchangeable sequences, general sufficiency is equivalent to predictive sufficiency (Bernardo and Smith, 2000). Because $p(x_{n+1}|x_{1:n})=g(f(x_{1:n}))\;\forall n$, $f(x_{1:n})$ is a predictive sufficient statistic for the sequence, and it is also a sufficient statistic. ∎
The resulting sufficient statistic also fully specifies the posterior on the parameters of the generating distribution, $p(\theta|x_{1:n})$. Sufficient statistics are easily identified for a wide range of distributions, including all exponential family distributions (Bernardo and Smith, 2000), and are easy to represent. Since the LM’s predictive distribution decomposes into the form above, this theoretical result gives us strong predictions about the contents of embeddings for models trained on exchangeable data.
3.3 Case 2: Latent state models
In a latent state model, each $x_{i}$ is generated based on a latent variable $z_{i}$. These $z_{i}$ are interdependent, with $z_{i}$ being generated from a distribution conditioned on $x_{i-1}$. Common latent state models include Kalman filters (where $x_{i}$ and $z_{i}$ are continuous and $p(x_{i}|z_{i})$ and $p(z_{i}|z_{i-1})$ are linear-Gaussian) (Kalman, 1960) and hidden Markov models (where the $z_{i}$ are discrete) (Baum and Petrie, 1966). In a latent state model, the posterior predictive distribution is
[…]
In this case, the posterior distribution $p(z_{n+1}|x_{1:n})$ captures all of the information in $x_{1:n}$ relevant to predicting $x_{n+1}$, rendering $x_{n+1}$ independent of $x_{1:n}$ when conditioned on the posterior, and acts as a predictive sufficient statistic in this model.
More formally, we can consider an embedding a representation $\phi_{n}$ such that a fixed operator $g(\phi_{n})$ can produce $p(x_{n+1}|x_{1:n})$. The posterior distribution $p(z_{n+1}|x_{1:n})$ satisfies this characterization, with Equation 2 showing that the relevant operator is the integral of $p(x_{n+1}|z_{n+1})$ over $z_{n+1}$. That operator can be easily approximated linearly and hence by a single layer of a neural network. The embedding $\phi_{n}$ thus need only represent $p(z_{n+1}|x_{1:n})$.
3.4 Case 3: Discrete hypothesis spaces
In a more specific version of Case 1, assume each $x_{i}$ is generated independently from some unknown generative model. Let $\mathcal{H}$ denote the set of hypotheses $h$ about the identity of this model. In this case, $x_{1:n}$ are exchangeable, but any sufficient statistics might be difficult to identify. The posterior predictive distribution can be written as
$p(x_{n+1}|x_{1:n})=\sum_{h\in\mathcal{H}}p(x_{n+1}|h)p(h|x_{1:n}).$
By an argument similar to Case 2, $p(h|x_{1:n})$ is a predictive sufficient statistic in this model and an embedding thus need only capture this posterior distribution.
3.5 Probing embeddings to recover predictive sufficient statistics
The three cases identified above specify information that should be encoded in the embeddings formed in neural networks. This sets up the second element of our approach, which is building probes to check whether this information is actually encoded in the representations of specific networks. We focus on transformers (Vaswani et al., 2017), since they are widely used as language models. We denote the target to be decoded for each sequence $x_{1:n}$ by a vector $t$. For example, in the discrete hypothesis space case the target $t$ is $p(h|x_{1:n})$, which is a vector on the simplex with vertices corresponding to different values of $h$. Exact choices of $t_{n}$ in each dataset are given in the next section. Given a trained transformer, we decode $t$ by by training a second model (a probe) to predict the target $t$ from the embedding $\phi_{n}$ of the corresponding document $x_{1:n}$. The probe $g$ maps from the document embedding to the target. To ensure that the relevant statistical information is contained in the LM, not in the probe, we keep the probe simple by defining it as a linear layer with softmax activations: $g(\phi_{n})=\text{Softmax}(\text{Linear}(\phi_{n}))$.
4 Empirical analysis of embeddings
We have identified three cases where a predictive sufficient statistic is expected to be contained in an autoregressive model’s embeddings. In this section, we conduct probing on transformers trained on different datasets to empirically verify this hypothesis. We also conduct ablation studies to further test this idea. In all experiments, we use a linear probe unless otherwise mentioned, three-layer transformer decoder, and the Adam optimizer (Kingma and Ba, 2014) (more implementational details in Appendix A.3).
4.1 Exchangeable models
4.1.1 Generative model
We evaluate our approach on three exchangeable models: the Gaussian-Gamma, Beta-Bernoulli, and Gamma-Exponential models. Due to space limitations we only describe the Gaussian-Gamma model in detail, but the full generative process underlying the other two models is given in Appendix A.1.
The Gaussian-Gamma model generates from a Gaussian distribution with two unknown parameters. For $M$ sequences $\{x^{(k)}\}_{1\leq k\leq M}$ with $N$ tokens each, each sequence is i.i.d. generated by a mean and precision parameter sampled from a prior. Across the $M$ sequences, $M$ mean and precision parameters are sampled to generate these sequences. For sequence ${x}^{(k)}$, the generative process is,
$\tau_{k}$ $\sim\textsc{Gamma}(\alpha_{0},\beta_{0})$
$\mu_{k}|\tau_{k}$ $\sim\mathcal{N}(\mu_{0},(\lambda_{0}\tau_{k})^{-1})$
$x^{(k)}_{n}|\mu_{k},\tau_{k}$ $\sim\mathcal{N}(\mu_{k},\tau_{k}^{-1})\;\text{for}\;n\in\{1,...,N\},$
where $\alpha_{0},\beta_{0},\mu_{0},\lambda_{0}$ are fixed hyperparameters. The predictive distribution for the next token in sequence ${x}$ is,
$p(x_{n+1}|x_{1:n})=\int p(x_{n+1}|\mu,\tau)p(\mu,\tau|x_{1:n})d(\mu,\tau).$
The optimal Bayesian agent uses the same prior distributions as the Gamma and Normal that generate $\mu$ and $\tau$. It will analytically infer the ground truth posterior $p(\mu,\tau|x_{1:n})$ for any stream of data it sees, and use this posterior for predicting next tokens. A Bayesian agent can also use other suitable priors and converge to the optimal posterior. To be consistent with other exchangeable conjugate models, we denote $\theta=(\mu,\tau)$ to indicate latent variables whose posterior distribution is predictive sufficient.
(a) Gaussian-Gamma.
(b) Beta-Bernoulli.
(c) Gamma-Exp.
Figure 2: Probe recovery of transformer-learned sufficient statistic (blue) and ground truth sufficient statistic (red), across 1000 test datapoints. The first row shows parameters probed in the non-OOD case (from left to right: Gaussian mean $\mu$, Gaussian precision $\tau$, Bernoulli mean, and Exponential mean). The second row shows the corresponding information in the OOD case.
4.1.2 Probing experiments
Experimental setup
We hypothesize that a transformer trained on this dataset should come to represent the sufficient statistics of the corresponding distribution. We use three probing setups.
Sufficient statistic. As suggested by the Proposition in Section 3.2, we probe the sufficient statistic for $\theta$. The probe uses the last token embedding of the transformer as input.
Moments of the posterior distribution. A consequence of knowing the sufficient statistic is finding the true posterior $p(\theta|x_{1:N})$, so we decode the moments of this posterior from the transformer embedding. The moments are functions of the mean and variance of the stream of data seen so far by the autoregressive predictor. In these conjugate models, it might be unsurprising for a transformer to encode the mean of the sequence it processes, because the optimal strategy for its loss function is to always predict the mean of the sequence that it sees so far. However, encoding the variance would not be directly related to this strategy and would support the argument that it infers sufficient statistics.
Out-of-distribution simulations. The analytical nature of the Bayesian predictor means that it is robust to datasets generated far from the prior – it would simply update its posterior based on the data. Thus, we probe the transformer on out-of-distribution (OOD) datasets that are generated from a distribution of the same form but with distinct hyperparameters.
Implementation details
For Gaussian-Gamma hyperparameters $\alpha_{0},\beta_{0},\mu_{0},\lambda_{0}$ we choose $\{5,1,1,1\}$. The OOD dataset is instead generated with hyperparameters $\{2,1,5,1\}$, so it is centered on a different mean (5 vs. 1) and has a higher spread compared to the original data. An analogous design was used for the other exchangeable conjugate models, with details in Appendix A.2.
(a) Gaussian-Gamma (linear probe).
(b) Gaussian-Gamma (two-layer probe).
Figure 3: Probe recovery of transformer-learned posterior distribution moments (blue) and ground truth moments (red) across 1000 test datapoints. The first row shows parameters probed in the general non-OOD case. The second row shows corresponding information in the OOD case.
Figure 4: Probing over the first 10 tokens themselves using the 10th token embedding of the transformer. Aside from perfectly encoding the 10th token, this embedding does not show memorization over the other 9 tokens as suggested by the noise in probe recovery.
Results
Results for all three models are shown in Figure 2, where the probe decodes the sufficient statistics. The remaining figures focus on the Gaussian-Gamma case (see Appendix A.3 for the other models). Figure 3(a) shows that the probe decodes the moments of the posterior distribution. Because the higher distribution moments are more volatile in value and less directly related to estimating the parameters $\theta$, we examine whether existing discrepancies are caused by an overly simple probe. We perform a second set of experiments where the probe has a hidden layer with ReLU activations (Figure 3(b)), showing stronger alignment.
Analyzing whether the transformer memorizes tokens
The transformer embedding (of size 128) can easily memorize a set of tokens, so we tested whether memorization is leading to the successful recovery of the sufficient statistics. We started by taking the 10th token embedding and probing the first 10 tokens’ sufficient statistics in each sequence, with high performance shown in Figure 8 in Appendix A.3. Then, we probed the token values themselves to look for memorization. Figure 4 suggests that memorization is generally absent. The 10th token embedding recovers the 10th token perfectly, but cannot recover the other 9 tokens. For the other 9 tokens, a correlation exists between probe results and true token values, but some level of correlation is expected because even seeing a single token can reveal information about this sequence’s generating distribution. However, the noise suggests that memorization is less of a goal than finding sufficient statistics.
4.2 Hidden Markov model
4.2.1 Generative model
For an HMM generated data with $M$ sequences with $N$ tokens each, we formulate the generative process as, for a sequence ${x}$,
$A_{c}$ $\sim\textsc{Dirichlet}_{C}(\gamma)\;\text{for}\;c\in\{1,...,C\}$
$B_{c}$ $\sim\textsc{Dirichlet}_{V}(\delta)\;\text{for}\;c\in\{1,...,C\}$
${z}_{0}$ $\sim\textsc{Categorical}({\pi})$
${x}_{i}$ $\sim\textsc{Categorical}(B_{{z}_{i}})$
${z}_{i+1}$ $\sim\textsc{Categorical}(A_{{z}_{i}}),$
where $C,V,{\pi}$ are initialized and denote, respectively, the number of classes, vocabulary size, and a list of probabilities on the number of classes to initialize the first latent state. $\gamma,\delta$ are scalar hyperparamters that are also initialized and fixed, and they represent the evenness of the samples from the Dirichlet distributions. $A$, $B$, and ${z}$, as a result, represent the transition matrix, the emission matrix, and the latent states, respectively.
We implement a forward-backward algorithm (Rabiner, 1989) to compute the posterior $p({z}_{n+1}|{x}_{1:n})$, and explore whether this distribution can be decoded from the transformer embedding on ${x}_{1:n}$.
4.2.2 Probing experiments
Like our experiments for exchangeable sequences, we use probes to explore whether an autoregressive transformer represents a predictive sufficient statistic for non-exchangeable sequences.
Decoding latent measures other than the predictive sufficient statistic
To further examine the role of the autoregressive objective in encoding a predictive sufficient statistic, we find other possible targets and test on whether they are more difficult to decode. We implement the Viterbi algorithm, which finds the most likely sequence of latent states given the whole observed sequence, and we use Viterbi’s output as our target. This is related to but distinct from the predictive sufficient statistic, $p(z_{n+1}|x_{1:n})$.
Implementational details
We choose $C=4,V=64,\gamma=0.5$, and set ${\pi}$ to be uniform in our experiments. We also vary $\delta$ to control the level of difficulty: how distinct is one class from another.
Results
Based on the autoregressive objective, we hypothesized that a transformer trained on HMM sequences should encode the posterior distribution on latent states $p({z}_{n+1}|{x}_{1:n})$. Table 1 suggests that this is the case, reflected in terms of both accuracy of decoding and error in reproducing this posterior distribution (as measured by squared loss). Furthermore, decoding performance is better for the posterior than for several related quantities. It is more difficult to decode other quantities using the same embedding, or even to decode any of these quantities using the embedding on one token ahead, ${x}_{n+1}$. Several other quantities exhibit high volatility across random seeds, since the transition matrix has a greater influence on results when the tokens conditioned in the target posterior and the token used in the embedding have a mismatch.
Table 1: Probing target quantities in HMM dataset using different transformer token embeddings..
$\delta=0.5$ $\delta=1$
Target Quantity Embedding Accuracy $\uparrow$ Squared Loss $\downarrow$ Accuracy $\uparrow$ Squared Loss $\downarrow$
$p({z}_{n+1}|{x}_{1:n})$ ${x}_{1:n}$ $90.8\pm 5.5\%$ $0.011\pm 0.012$ $90.4\pm 5.9\%$ $0.011\pm 0.01$
$p({z}_{n+1}|{x}_{1:n})$ ${x}_{1:n+1}$ $66.6\pm 20.9\%$ $0.072\pm 0.042$ $65.9\%\pm 21.6\%$ $0.058\pm 0.035$
$p({z}_{n+1}|{x}_{1:n+1})$ ${x}_{1:n}$ $53.2\pm 11.2\%$ $0.356\pm 0.043$ $50.8\pm 11.3\%$ $0.278\pm 0.038$
$p({z}_{n+1}|{x}_{1:n+1})$ ${x}_{1:n+1}$ $86.5\pm 4\%$ $0.066\pm 0.034$ $82.2\pm 5.8\%$ $0.067\pm 0.036$
$\hat{{z}}_{n+1}$ ${x}_{1:n}$ $59.8\%\pm 12.1\%$ / $61.4\pm 14.7\%$ /
$\hat{{z}}_{n+1}$ ${x}_{1:n+1}$ $80.8\%\pm 7\%$ / $77.5\pm 8.8\%$ /
Note: The target $p({z}_{n+1}|{x}_{1:n})$ is a simplex vector found by running the forward-backward algorithm, and the target $\hat{{z}}_{n+1}$ is a scalar standing for the most likely latent class found by Viterbi algorithm. Average and standard deviation across 10 random seeds are reported.
4.3 Discrete hypothesis spaces
4.3.1 Generative model
We simulate a dataset where the data consist of a sequence of two-dimensional points uniformly sampled from a rectangular region in 2D space (cf. Tenenbaum, 1998). In this case, the hypothesis space $\mathcal{H}$ is defined as the set of all rectangles whose corner points are pairs of integers in $\{0,1,2,...,7\}$. For each sequence ${x}$, the generative process is,
$h_{\text{rect}}$ $\sim\mathcal{H},$
$x_{i}$ $\sim\textsc{Uniform}(h_{\text{rect}})\;\forall i.$
The generative process uniformly samples a rectangle $h_{\text{rect}}$ from the set of all rectangles $\mathcal{H}$. Then, each token in a sequence is sampled uniformly from the region defined by $h_{\text{rect}}$.
(a) A hypothesis space with equal-width rectangles.
(b) A hypothesis space with unequal-width rectangles.
Figure 5: Two discrete hypothesis spaces $\mathcal{H}$ used in our experiments. Any continuous rectangle contained within the axes (e.g., the red or the orange rectangle) is a valid hypothesis $h\in\mathcal{H}$. This results in 784 distinct hypotheses in each hypothesis space. The data consist of a sequence of points sampled uniformly from the target rectangle.
4.3.2 Probing experiments
Implementational details
We use rectangles with seven unit blocks on each side, resulting in a size-784 hypothesis space, i.e., 784 different possible rectangles from which the sequences are drawn. The unit lengths are either 1 (Figure 5(a)), or 0.4 and 0.16 alternating (Figure 5(b)).
Probing results
We probe for the distribution $p(h|x_{1:n})$, which is a length-784 simplex vector, using the last token embedding of each sequence. Results are shown in Table 2. In general, the true hypothesis can be found with high accuracy, even though the number of classes is high.
Embedding visualization
We visualize the transformer embeddings to show what they represent in terms of this discrete set of hypotheses (Figure 6). We use Principal Components Analysis (PCA) to reduce the 128-dimensional embeddings to 2D, and color the points by properties of each rectangle. Embeddings are clustered into different regions based on the mean of the center of the true generating rectangle (Figure 6(a)), where, for instance, the top right yellow cluster corresponds to the upper right-most rectangle, and the middle teal cluster corresponds roughly to the many rectangles whose center have a mean of 3.5. Additionally, the embeddings encode the position along individual axes (Figure 6(b)), as well as distance between corners of the generating rectangle (Figure 6(c)).
(a) Colored by mean of the generating rectangle.
(b) Colored by mean along the y-axis of the generating rectangle.
(c) Colored by spread along the y-axis of the generating rectangle.
Figure 6: Two-dimensional representation of embeddings of all validation datapoints in the discrete hypothesis space dataset (the setup is unequal width and sample size $=50$). Subfigures show the same embeddings, colored by different properties of the true rectangle used to generate the data.
Table 2: Probing results on discrete hypothesis spaces. In general, the probe achieves strong performance in recovering a 784-length vector, and performance increases as task difficulty decreases. Average and standard deviation across 10 random seeds are reported.
Equal Width Unequal Width
Sample Size Accuracy $\uparrow$ Squared Loss $\downarrow$ Accuracy $\uparrow$ Squared Loss $\downarrow$
20 $87.3\pm 1.2\%$ $0.173\pm 0.016$ $66\pm 1.7\%$ $0.29\pm 0.012$
50 $99.5\pm 0.2\%$ $0.008\pm 0.004$ $88.5\pm 0.9\%$ $0.159\pm 0.011$
5 Discussion
Limitations and Future Directions
Several limitations of our analysis motivate future work.
Extending to more complex Bayesian probabilistic models. We focused on three cases that consist of a one-step generation process and a simple likelihood $p(x|z)$ that can be approximated by a linear layer. Our analysis can be extended to hierarchical models or likelihoods parameterized by, for instance, neural networks, while simultaneously exploring more choices of transformer embeddings.
Interpreting neural networks with Bayesian probabilistic models. A similar Bayesian approach could be used to interpret the representations learned by other autoregressive neural networks. A direction for future work is extending our analyses to deep learning models for other modalities.
Task difficulty. As seen in the hypothesis space experiments, the ability to decode predictive sufficient statistics decreases as task difficulty increases. Developing a better understanding of the relationship between task difficulty and transformer embeddings is an important direction for future work.
Conclusion
We have developed a general framework for analyzing what the embeddings of an autoregressive language model should represent. Our analyses suggest that such embeddings should represent latent structures such that the next token $x_{n+1}$ is independent from previous tokens $x_{1:n}$ when conditioned on that structure, a property possessed by predictive sufficient statistics. We confirmed this hypothesis with probing experiments on three cases where predictive sufficient statistics can be identified. We hope that our findings contribute to bridging the gap between Bayesian probabilistic models and deep neural networks.
Acknowledgements
This work was supported by ONR grant number N00014-23-1-2510.
Appendix A Appendix
A.1 Definition of additional exchangeable conjugate models
We consider the generative process for sequence $x_{i}$, where $x_{ij}$ are i.i.d. across $j$.
A.1.1 Beta-Bernoulli model
$\theta_{i}$ $\sim\textsc{Beta}(\alpha,\beta)$
$x_{ij}$ $\sim\textsc{Bernoulli}(\theta_{i}),$
where $\alpha,\beta$ are fixed hyperparameters.
A.1.2 Gamma-Exponential model
$\theta_{i}$ $\sim\textsc{Gamma}(\alpha,\beta)$
$x_{ij}$ $\sim\textsc{Exponential}(\theta_{i}),$
where $\alpha,\beta$ are fixed hyperparameters.
A.2 Implementational details
All training processes are run on a T4 or A100 Nvidia GPU.
Experimental process
Each dataset is split into three sets: set 1, set 2, and set 3. Set 1 is used for training the transformer. Set 2 is used for validating the transformer and getting embeddings from transformer that are used to train the probe. Set 3 is used for validating the probe.
With the exception of the discrete hypothesis space datasets, the sizes for the three sets are: 10000, 3000, 1000, and each sequence is 500-tokens long. In the discrete hypothesis space datasets, we experimented with different sequence lengths (detailed in our results), and the sizes for the three sets are: 20000, 19000, 1000.
Transformer
We use a three-layer transformer decoder with hidden-size $=128$ and number of attention heads $=8$. If the input is categorical (similar to tokens in natural corpus), we employ the standard word embedder layer before the decoder layers. If the input is continuous, we use a Linear layer to map inputs to dimension 128 in place of the word embedder layer.
Dropout $=0.1$ is applied, and learning rate $=0.001$, batch-size $=64$.
Probe
The probe is a linear layer with softmax activations. Learning rate is tuned in $[0.001,0.01]$, and batch-size $=64$.
Hyperparameters for exchangeable conjugate models
On the Beta-Bernoulli model, we use $\alpha=2,\beta=8$. In the OOD case, $\alpha=8,\beta=2$.
On the Gamma-Exponential model, we use $\alpha=2,\beta=4$. In the OOD case, $\alpha=2,\beta=1$.
A.3 Experimental results
Figure 7 shows results on posterior distribution moments on Beta-Bernoulli and Gamma-Exponential models. On the Gamma-Exponential model, we divide the target second, third, and fourth moments by factors of 10, 100, 1000, respectively so that each moment is given roughly equal importance.
(a) Beta-Bernoulli.
(b) Gamma-Exponential (linear probe).
(c) Gamma-Exponential (two-layer probe).
Figure 7: Probe recovery of transformer-learned posterior distribution moments (blue) plotted with ground truth moments (red). The first row shows parameters probed on 1000 test datapoints in the general, i.e., non-OOD, case. The second row shows corresponding information in the OOD case.
Figure 8 shows Gaussian-Gamma probing results on sufficient statistics on the 10th token, complementing our exploration of token memorization in transformer (Figure 4).
Figure 8: Probing over mean (left) and standard deviation (right) of the first 10 tokens using the 10th token embedding of the transformer in the Gaussian-Gamma dataset. The first row corresponds to the same generation process, and the second row corresponds to the OOD case.